The decimal equivalent of the binary number (11010.101)2
is
(A) 9.625 (B) 25.265 (C) 26.625 (D) 26.265

Grade:

1 Answers

Arun
3 years ago
The subscript indicates base :
(11010.101)= 1 x 2+ 1 x 2+ 0 x 2+ 1 x 2+ 0 x 2​0 + 1 x 2-+ 0 x 2-+ 1 x 2-3 =
16+8+0+2+0+1/2+0+1/8 = 26+0.625 = 26.625